[ Bruce is 36 years old. His son is 8. In how many years will Bruce be 3 times as old as his son? ]
Answer by TimothyLamb(4379)

---
b = 36
s = 8
36 + x = 3(8 + x)
36 + x = 24 + 3x
2x = 12
x = 6
---
6 years from now
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php